OpenStudy (anonymous):

Simplify the expression below. 4x(3x - 7) - 19x2 Select one of the options below as your answer:A. -35x2B. -31x2 - 28xC. -7x2 - 28xD. -7x2 - 7

OpenStudy (anonymous):

Okay so we have 4 * 3 = 12 and x*x = x^2, so 12x^2 - 28x (because 4*7 is 28 and there is only one x so tack it on at the end)..12x^2 - 28x - 19x^2. Combine your like terms, -7x^2 - 28x..so answer is C assuming they don't want you to re-factor it

OpenStudy (anonymous):

ohkay thank you

OpenStudy (anonymous):

welcome